OpenStudy (anonymous):

What is the next number in the pattern? 1,-4,9,-16...

jimthompson5910 (jim_thompson5910):

Notice how 1, 4, 9, 16 is the list of the first four perfect squares 1 = 1^2 4 = 2^2 9 = 3^2 16 = 4^2

jimthompson5910 (jim_thompson5910):

the only difference is that you are alternating in sign for each term

OpenStudy (anonymous):

so it grows by multiplication?

jimthompson5910 (jim_thompson5910):

the rule for generating perfect squares is n^2 where n is a positive whole number the rule for your sequence is (-1)^(n+1)*n^2, again n is a positive whole number the (-1)^(n+1) portion helps you get the correct sign for each term

OpenStudy (anonymous):

i dont really understand.... im sorry

jimthompson5910 (jim_thompson5910):

have you learned about perfect squares?

OpenStudy (anonymous):

no not really i had a virtual class and they didnt really go into that topic

jimthompson5910 (jim_thompson5910):

have you learned about squaring numbers?

OpenStudy (anonymous):

ummm... can you give me an example of squaring numbers? it sounds familiar

jimthompson5910 (jim_thompson5910):

squaring means you multiply a number by itself

jimthompson5910 (jim_thompson5910):

it comes from the idea that the area of a square is (side)*(side)

jimthompson5910 (jim_thompson5910):

example 9^2 = 9*9 = 81 area of a square with side length of 9 is 9*9 = 81 

OpenStudy (anonymous):

yes i understand that and i have learned it

jimthompson5910 (jim_thompson5910):

so what is 5^2

OpenStudy (anonymous):

25

jimthompson5910 (jim_thompson5910):

how about 6^2

OpenStudy (anonymous):

36

jimthompson5910 (jim_thompson5910):

so, 1 = 1^2 4 = 2^2 9 = 3^2 16 = 4^2 agreed?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

1,4,9,16 are a list of perfect squares they are the first 4 perfect squares (squaring the first four positive whole numbers)

jimthompson5910 (jim_thompson5910):

the next perfect square is 25 since 5^2 = 25

jimthompson5910 (jim_thompson5910):

we have this list now: 1, 4, 9, 16, 25

jimthompson5910 (jim_thompson5910):

let's go back to 1,-4,9,-16...

jimthompson5910 (jim_thompson5910):

the difference between 1,-4,9,-16... and 1, 4, 9, 16,... is that the signs are alternating

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

so the next term is 25 1,-4,9,-16, 25 and the next term is -36 1,-4,9,-16, 25, -36 etc etc

OpenStudy (anonymous):

yes it makes sense thank you very much

jimthompson5910 (jim_thompson5910):

you're welcome